OpenStudy (anonymous):

Solve system using substitution 3x+y=5 2x-5y=9

OpenStudy (amistre64):

just for linear practice :) 3x+y=5 2x-5y=9 x = (9+25)/(2+15) y = (27-10)/(-15-2) maybe?

OpenStudy (anonymous):

solve for one variable first: i'm going to use the first equation: 3x+y=5 so y=-3x+5 plug that in to equation 2 2x-5(-3x+5)=9 2x+15x-25=9 17x-25=9 17x=34 x=2 plug that in: 3(2)+y=5 6+y=5 y= -1 so if all is correct, x=2, y=-1

OpenStudy (amistre64):

trying to do a determinant backwards is tough

OpenStudy (amistre64):

x = (9+25)/(2+15) = 34/17 = 2 y = (27-10)/(-15-2) = 17/-17 = -1 id say its right ;)

OpenStudy (anonymous):

Thank you all! Life saved.